OpenStudy (anonymous):

What is the square root of 3+3/3+12?

OpenStudy (anonymous):

Is that the correct order of operations? Or do you need to throw in some parenthesis to clarify the situation?

OpenStudy (anonymous):

I made the question just like it looks. I know the answer (I made up the problem), I'm just bored and wanna see if anyone else can figure it out. :P

OpenStudy (anonymous):

As it is, I would say you have sqrt(3 + 3/3 + 12) = sqrt(3 + 1 + 12) = sqrt(16) = 4

OpenStudy (anonymous):

Close.

OpenStudy (anonymous):

+/- 4

OpenStudy (anonymous):

Haha you tricksy goon you.

OpenStudy (anonymous):

Ah... I was hoping it'd take longer for someone to figure it out. Lolz. Well played.

OpenStudy (anonymous):

I'm a future math teacher. We have to know all the tricks.

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

Ah... a great career. I would have loved to do it, but the kids here who just want answers given to them instead of learning the magic behind math make me really sad. That, and I hear there's not a lot of money in it. :(

OpenStudy (anonymous):

Well, yeah. That's certainly a problem on here. But there are people on here who want to understand the math. I just try to be selective about who I help. If they get impatient with me for not handing them an answer, I move on. Unfortunately the problem is somewhat caused by the approach taken in many math classrooms. This constant idea that as long as the number in the box at the bottom of the page is correct, that's all that matters. Teachers have to ask more of their students. Demand that they understand the math and be able to use it to solve real problems. Aaand of course there's not much money in teaching, but I'm excited about it anyhow. I think it'll be worth it.